OpenStudy (anonymous):

Chase wants to factor x2 + 10x + 25 by grouping; however, Paige says it is a special product and can factor a different way. Using complete sentences, explain and demonstrate how both methods will result in the same factors.

OpenStudy (anonymous):

@Amandakay1919

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Hero

OpenStudy (the_fizicx99):

IT has an "x^2" so in both parenthesis it'll have: (x )(x ) Now find 2 numbers that when multiplied gives you 25 but when added gives you 10, 5 and 5, 5*5 = 25, 5 + 5 = 10 (x + 5)(x + 5) The other one is, it's a perfect square trinomial. You could simplify it down to (x + 5) (x+5)^2 = x^2 + 2(5)(x) + 5^2 = x^2 + 10x + 25

OpenStudy (anonymous):

thanks can you help me with another 1 of my quesations

OpenStudy (the_fizicx99):

o.0 Questions? Erm Can't I gtg :/ @Luigi0210 <-- my sister can help you :>